### v1

# Action Spotting and Precise Event Detection in Sports: Datasets, Methods, and Challenges

Abstract
Video event detection has become an essential component of sports analytics, enabling automated identification of key moments and enhancing performance analysis, viewer engagement, and broadcast efficiency. Recent advancements in deep learning, particularly Convolutional Neural Networks (CNNs) and Transformers, have significantly improved accuracy and efficiency in Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES). This survey provides a comprehensive overview of these three key tasks, emphasizing their differences, applications, and the evolution of methodological approaches. We thoroughly review and categorize existing datasets and evaluation metrics specifically tailored for sports contexts, highlighting the strengths and limitations of each. Furthermore, we analyze state-of-the-art techniques, including multi-modal approaches that integrate audio and visual information, methods utilizing self-supervised learning and knowledge distillation, and approaches aimed at generalizing across multiple sports. Finally, we discuss critical open challenges and outline promising research directions toward developing more generalized, efficient, and robust event detection frameworks applicable to diverse sports. This survey serves as a foundation for future research on efficient, generalizable, and multi-modal sports event detection.
1 Introduction
Sports represent one of the largest global markets, projected to reach 599.9 billion US dollars by 2025 and 826 billion US dollars by 2030, with a compound annual growth rate of 6.6% [1]. Beyond its economic impact across industries such as media, marketing, and apparel, sports fundamentally center around athletic performance, where optimizing player efficiency, refining game strategies, and enhancing fan engagement are critical.
The rise of sports analytics—the systematic collection, processing, and analysis of performance data—has enabled data-driven decision-making, leading to fundamental shifts in strategy. For instance, basketball has seen an increased reliance on three-point shooting, guided by predictive models that estimate expected points from various court locations [2].
Within this context, video event detection has emerged as a fundamental yet challenging task in sports analytics. Accurate identification of key moments—such as corner kicks in soccer, rally conclusions in racket sports, or scoring events across disciplines—provides critical insights for coaches and athletes, supporting more effective performance analysis and tactical decision-making. Moreover, event detection benefits downstream applications by filtering non-playing segments, optimizing computational resources for subsequent tasks such as object tracking, and enabling automated highlight generation for commercial broadcasting and fan engagement.
Computer Vision (CV) has played a pivotal role in advancing sports video analysis, enabling automated player tracking, action recognition, tactical analysis, injury prevention, and officiating through advanced visual data processing [3, 4, 5, 6, 7, 8]. While traditional CV methods laid the foundation for many applications [5, 9], they heavily relied on handcrafted features and struggled with dynamic environments, occlusions, and real-time constraints.
Recent breakthroughs in deep learning (DL), particularly through Convolutional Neural Networks (CNNs) and Transformer architectures, have significantly enhanced the accuracy, efficiency, and scalability of sports video analysis tasks [5]. These modern techniques now power real-time object detection, robust player tracking, pose estimation, and fine-grained event recognition, fundamentally reshaping the landscape of sports analytics.
This survey provides a comprehensive analysis of DL-based CV techniques for video event detection in sports—a field that has seen remarkable advancements with the emergence of modern neural architectures and the formulation of distinct problem settings, notably Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES).
Table 1: Comparison of Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES).
Aspect TAL AS PES
Output Type Temporal interval Single key frame Single key frame
Annotation Format Start and end times Single timestamp Single timestamp
Tolerance Window $\sim$1–5 seconds 5–60 frames 0–2 frames
Best Suited For Long-duration actions Ambiguous, fast-paced actions Frame-accurate event detection
Annotation Cost High Medium High
Use Cases Generic action localization Sports highlight detection Fine-grained critical events
Previous studies, such as by Ghosh et al. [5], broadly survey AI applications in sports analytics but do not specifically address DL-based CV methods. Similarly, Thomas et al. [9] focus primarily on traditional CV approaches designed for multi-camera systems. In contrast, our survey specifically targets deep learning-based approaches—including recent advances in CNNs and Vision Transformers—for event detection tasks within monocular video contexts, enhancing its relevance for real-world deployment.
Other comprehensive surveys by Naik et al. [3], Zhao et al. [4], Kamble et al. [10], and Wu et al. [6] extensively review computer vision methods across various sports analytics tasks, including object tracking and action recognition. However, they do not specifically address video event detection. The most closely related work by Hu et al. [11] provides an extensive overview of TAL methods, but it does not cover the increasingly critical tasks of AS and PES, nor is it specifically focused on the sports domain.
This survey addresses these significant gaps by making the following key contributions:
  • Explicit Definition and Analysis of AS and PES: We clearly define, differentiate, and critically review AS and PES tasks, highlighting their significance and unique challenges within sports video analytics.
  • Comprehensive Methodological Survey: We systematically review DL-based approaches for sports event detection, including techniques based on convolutional networks, transformer architectures, and multi-modal fusion.
  • Evaluation Metrics and Datasets Review: We provide a thorough summary of evaluation protocols and benchmark datasets relevant to TAL, AS, and PES tasks.
  • Identification of Open Challenges: We discuss key limitations of existing methods, including issues related to temporal precision, multi-sport generalizability, and data efficiency, and outline promising research directions.
By clearly delineating these specific and timely contributions, our survey offers a distinct and valuable perspective within the broader sports analytics literature.
Figure 1: Example of Temporal Action Localization: in tennis, the time interval of the whole serve action has been labeled (blue bar).
Figure 2: Example of Precise Event Spotting: in table tennis, the moment a player contacts the ball during a serve (red) or when a ball bounces on the table (blue).
2 Taxonomy of Sports Event Detection
In the domain of sports video analysis, three primary tasks relevant to event detection have been defined: TAL, AS, and PES. While these tasks share certain similarities, each addresses distinct objectives and annotation requirements, as summarized in Table 1. The key differences among these tasks are discussed in detail in the following subsections.
2.1 Temporal Action Localization
Temporal Action Proposal Generation (TAPG) aims to identify temporal segments within videos that are likely to contain actions or events, without necessarily assigning specific action labels. Building upon TAPG, Temporal Action Localization (TAL)—also referred to as Temporal Action Detection (TAD)—extends the task by not only detecting these temporal segments but also classifying the actions occurring within them [12, 13, 11], as illustrated in Figure 1.
Among these tasks, TAPG is generally regarded as the most challenging, as accurately estimating the temporal boundaries of actions is inherently difficult. In contrast, once reliable proposals are available, action classification models tend to achieve relatively high accuracy [11]. This makes TAL particularly effective for recognizing everyday actions that unfold over extended periods.
2.2 Action Spotting
Introduced by Giancola et al. in 2018 [14], AS—also referred to as Event Spotting—addresses the ambiguity inherent in defining precise temporal boundaries for actions. Several factors contribute to this challenge: some actions occur extremely rapidly, others have continuous and indistinct endpoints, and many actions may overlap or conflict with one another. These ambiguities are especially pronounced in sports scenarios. For instance, in soccer, accurately determining the start and end times of events such as ”scoring a goal” or ”passing” is often nontrivial, due to the continuous nature of gameplay and the lack of sharply defined action boundaries.
To address these challenges, AS proposes an alternative to TAL by representing each action with a single keyframe rather than a temporal interval. While both TAL and AS aim to temporally locate actions within untrimmed videos, they differ fundamentally in annotation granularity. TAL annotates actions with explicit start and end times, making it suitable for longer-duration activities. In contrast, AS captures brief and rapidly occurring events using a single timestamp, thereby simplifying annotation efforts and mitigating ambiguities associated with temporal boundaries.
2.3 Precise Event Spotting
While AS facilitates coarser detection by tolerating errors within relatively broad temporal windows (typically between 5 and 60 seconds), certain applications in sports demand much higher temporal precision. Hong et al. [15] highlighted that in sports such as tennis [16] and figure skating [17], even a margin of error of 1–2 frames can result in missing critical events, such as the precise moment of ball contact in tennis.
Similarly, consultations with Table Tennis Australia confirmed that frame-level precision is critical in table tennis, where even minor inaccuracies can significantly impact the determination of ball bounce locations, given the rapid motion of the ball across frames in professional-level matches, as illustrated in Figure 2.
To meet these demands, PES was proposed as an extension of AS. PES requires extremely strict detection thresholds, typically constrained within 1–5 frames (approximately 33–200 milliseconds at standard frame rates of 25–30 FPS). PES thus ensures highly accurate event localization at the frame level, addressing the critical need for fine-grained temporal precision in professional sports analysis.
3 Video Event Detection
Generally, a video event detection model do this task in two step:
  • Temporal Segmentation: This involves identifying the timing of an event, either by locating its start and end timestamps within a continuous video (as in Temporal Action Localization, TAL) or by detecting a single representative keyframe (as in Action Spotting, AS, and Precise Event Spotting, PES)
  • Event Classification: Once an event is segmented, the model classifies it into predefined categories, such as team events (e.g., goals, time-outs) or individual actions (e.g., shots, fouls, serves).
In this section, we dive into the review of video event detection.
3.1 Temporal Action Localization
Current methods for TAPG typically fall into two main categories: Global-to-Local and Local-to-Global approaches. Global-to-Local methods generate action proposals through sliding windows or pre-defined anchor structures, segmenting videos into fixed-length or hierarchically structured intervals to localize actions. Local-to-Global methods independently evaluate each frame to predict the likelihood of action start, action end, and overall actionness. Proposals are subsequently created by connecting high-confidence start and end frames, followed by a confidence-based evaluation to confirm whether the proposal indeed contains an action. The difference between these two approaches is shown in figure 3. The subsequent sections will explore these two approaches comprehensively, discussing their strengths, limitations, and practical applications in sports event detection.
Figure 3: Comparison of Global-to-Local (GTL, left) and Local-to-Global (LTG, right) approaches. In GTL, predefined temporal anchors are classified as action or background, followed by a regressor that refines the event’s time interval. In contrast, LTG first predicts per-frame start and end probabilities to define the temporal boundaries and then classifies the resulting interval.
J. Gao et al. proposed the Temporal Unit Regression Network (TURN) to address the temporal action proposal generation problem [18]. TURN first decomposes an untrimmed video into short video units, which serve as fundamental processing blocks, with visual features extracted from each unit using a base model C3D [19]. To construct meaningful clip-level representations, TURN pools a set of contiguous units across multiple temporal scales, forming a clip pyramid. Each clip is further enriched by concatenating features from both its internal and surrounding units, ensuring contextual information is preserved. At this stage, each clip is treated as a proposal candidate, and TURN outputs a confidence score indicating whether the clip contains an action instance, along with two regression offsets that refine the predicted start and end times of the action within the clip. Finally, Non-Maximum Suppression (NMS) is applied to eliminate redundant proposals, improving localization accuracy and ensuring high-quality action boundaries. They have also proposed a new metric for this task which is called Average Recall vs. Frequency of retrieved proposals (AR-F) which covers two criteria, (1) it should be capable of evaluating the performance of different methods on the same dataset efficiently; (2) it should be capable of evaluating the performance of the same method across different dataset. Even though TURN has achieved the SOTA performance on two challenging datasets THUMOS-14 [20] and ActivityNet [21], with the highest AR at the time because of the sliding window nature where it covers all segments in the video however the temporal boundaries are not precise enough, indicating its flaws when accurately deciding the correct time frame for the event [22, 13, 23].
On the other hand, Temporal Action Grouping (TAG) [24] was proposed as an alternative to fixed temporal scale methods, drawing inspiration from the watershed algorithm [25]. TAG groups continuous high-score regions into proposals rather than relying on predefined temporal scales. It applies a binary classification to each unit or snippet (a few contiguous frames) to generate actionness scores, which indicate the likelihood of an action occurring in that unit. This method is capable of producing high-precision temporal boundaries, provided that the actionness scores are accurate. However, its performance heavily depends on the quality of the predicted actionness scores, making it less effective when predictions are unreliable [22].
To address these limitations, Complementary Temporal Action Proposal (CTAP) was introduced by J. Gao et al. [22], leveraging the strengths of both actionness-based proposals and sliding-window proposals. CTAP first generates both types of proposals and then applies a proposal complementary filter, which collects missing correct proposals that TAG alone might overlook by retrieving them from the sliding-window proposals. In this stage, a binary classifier is used, taking a sequence of unit features as input and outputting a probability score that indicates whether a proposal can be accurately detected by the unit-level actionness model and TAG. Finally, a Temporal Convolutional Adjustment and Ranking Network is employed to rank proposals and refine their boundaries. Experimental results on THUMOS-14 and ActivityNet v1.3 demonstrate that CTAP outperforms TURN in AR, highlighting its effectiveness in improving proposal quality. However, the combination of multiple proposal generation methods can impact computational efficiency, potentially increasing processing time.
Figure 4: Examples of events defined in the context of soccer. From top to bottom: Goal—the instant the ball crosses the goal line; Substitution—the moment a player enters the field; Card—the instant the referee shows a card to a player. All examples are from SoccerNet [14].
T. Lin et al. proposed the Boundary Sensitive Network (BSN), the first approach designed to predict the start, end, and actionness of each frame in a video [12]. Proposals are generated by pairing start and end points with high probabilities, while low-confidence candidates are discarded by an evaluation module. BSN significantly outperforms previous methods and achieves high performance even with relatively few proposals. On the THUMOS14 dataset, BSN surpasses all SOTA methods across a range of proposal counts (10 to 1000). Notably, when the average number of proposals is 50, BSN improves AR from 21.86% to 37.46%, demonstrating its effectiveness in capturing action boundaries with fewer proposals.
However, BSN suffers from inefficiencies due to its proposal feature construction and confidence evaluation processes, which are applied individually to each proposal, leading to high computational overhead. To address this issue, Boundary-Matching Network (BMN) was introduced [13], incorporating a Boundary-Matching (BM) mechanism for more efficient confidence evaluation of densely distributed proposals. Instead of evaluating each proposal separately, BMN represents proposals as matching pairs of their start and end boundaries and constructs a two-dimensional BM confidence map. This map models densely distributed proposals by encoding continuous starting boundaries and temporal durations, allowing confidence scores to be computed simultaneously for all proposals. By leveraging this structure, BMN significantly improves efficiency while maintaining strong performance in temporal action proposal generation.
Building upon these advancements, BSN++ was introduced as an enhanced version of BSN to further refine proposal generation and confidence evaluation [23]. While BSN effectively captures action boundaries, it lacks a mechanism to assess the completeness of proposals, often leading to inaccurate segmentations. BSN++ addresses this by incorporating a proposal completeness modeling module, which evaluates whether a detected segment fully encapsulates an action instance. Additionally, BSN++ enhances feature representation by introducing context-aware proposal features, ensuring that both internal and surrounding contextual information contribute to more precise action localization. To improve proposal ranking, a global-local confidence fusion module is integrated, leveraging both local proposal confidence and global consistency to reduce false positives. Furthermore, BSN++ adopts the BM confidence map to evaluate multiple proposals simultaneously, significantly improving computational efficiency. These refinements enable BSN++ to generate higher-quality proposals while reducing redundancy, leading to superior performance in temporal action proposal generation.
However, Z. Qing et al. argue that existing methods struggle to effectively model temporal relationships, as neither convolution-based approaches nor global fusion mechanisms can fully capture long-range dependencies. Additionally, relying solely on internal or boundary context for regression is insufficient for generating proposals with precise boundaries [26]. To address these limitations, they proposed the Temporal Context Aggregation Network (TCANet) for high-quality proposal generation, which consists of two key components. First, the Local-Global Temporal Encoder (LGTE) is introduced to capture both local and global temporal relationships in a channel-grouping fashion. The Local Temporal Encoder (LTE) is responsible for processing the first A groups of the input feature maps, which are linearly transformed and evenly divided into N groups along the channel dimension to enhance short-term temporal modeling. Simultaneously, the remaining N - A groups are handled by the Global Temporal Encoder (GTE), which captures long-range dependencies. By combining these two encoders, LGTE effectively integrates long-term contextual information while preserving structural and fine-grained local details, leading to richer proposal representations. Second, the Temporal Boundary Regressor (TBR) is designed to refine proposal boundaries by leveraging both internal and boundary context. This is achieved through a dual regression approach: (1) Frame-level boundary regression, which refines the starting and ending locations of proposals with boundary sensitivity, and (2) Segment-level boundary regression, which adjusts the center location and duration of proposals based on a holistic perception of their structure. The final predicted proposal boundaries are obtained by fusing the outputs of both regression modules, ensuring precise action localization. Experimental results demonstrate that TCANet significantly enhances action proposal generation and action detection, outperforming existing methods by improving both boundary accuracy and proposal completeness.
Previous works typically use tIoU between the proposal and the ground truth action instance—referred to as the action score—to evaluate proposal confidence and develop high-quality temporal action proposals. However, H. Yang et al. argue that background information is also significant but was previously overlooked. Motivated by this observation, they proposed the Background Constraint Network (BCNet) to generate high-quality temporal action proposals [27]. BCNet consists of two main modules: the Action-Background Interaction (ABI) Module and the Boundary Prediction (BP) Module. Inspired by the Transformer [28], the ABI module first employs self-attention to capture relationships between frames and enhance feature representations. The enhanced frame features are then processed by the Difference-Attention Unit, which aggregates features based on the inconsistency between action and background features under supervision. Finally, a prediction head encodes both action and background features, followed by a Sigmoid layer to generate action and background probability sequences. During training, unlike previous works that rely solely on action score loss, BCNet incorporates background score loss to further refine proposal confidence. Extensive experiments on the ActivityNet-v1.3 and THUMOS14 datasets demonstrate the effectiveness of explicitly modeling background information and leveraging attention mechanisms in TAPG.
Table 2: Summary of methods for Action Spotting (AS) and Precise Event Spotting (PES). Performance is grouped under the Test and Challenge sets. All results are reported on the SoccerNet Action Spotting dataset, where italicized entries indicate results from SoccerNet-v1 [14], and plain text entries are from SoccerNet-v2 [29]. Bold numbers indicate the highest scores in each column. A ✓in the generalizability column means the method was tested on other sports beyond soccer.
Method Year Technique Modality Test Set Challenge Set Generalizability
Tight Loose Tight Loose
Giancola et al. [14] 2018 Temporal Pooling Visual 31.37 30.74
Rongved et al. [30] 2020 3D Convolution Visual 56.3
Vats et al. [31] 2020 Multi-Temporal Scales Visual 60.1
CALF [32] 2020 Context-aware Loss Visual 41.61 42.22
Vanderplaetsen et al. [33] 2020 Multi-Modal Fusion Visual + Audio 39.90
NetVLAD++ [34] 2021 Temporally Aware Pooling Visual 53.40 52.54
RMS-Net [35] 2021 Regression & Masking Visual 63.49 60.92
Zhou [36] 2021 Transfer Learning Visual 47.05 74.77 49.56 74.84
E2E-Spot [15] 2022 GRU + Gated Shift Module Visual 66.73 73.62
Shi et al. [37] 2022 Multiple Scene Encoders Visual 55.20
STE [38] 2022 1D Convolution Visual 58.29 71.58 58.71 70.49
SpotFormer [39] 2022 Multi Feature Extractors Visual 60.90 81.50
Soares et al. [40] 2022 Dense Detection Anchor Visual 65.07 78.59 68.33 78.06
Zhu et al. [41] 2022 Multi-Scale Vision Transformer Visual 52.04 60.86
ASTRA [42] 2023 Transformer Encoder-Decoder Visual + Audio 70.10 79.21
T-DEED [43] 2024 Scalable-Granularity Perception Visual
COMEDIAN [44] 2024 Self-Supervised Learning & Knowledge Distillation Visual 73.10 68.38 73.98
Tran et al. [45] 2024 Vision-Language Object Detector Visual + Text 62.49 73.98 69.38 76.15
3.2 Action Spotting & Precise Event Spotting
In contrast to TAL, which focuses on predicting temporal intervals for actions, AS and PES have gained significant attention in the sports analytics domain. AS and PES aim to identify actions or events using a single key frame, making them particularly suited for rapidly occurring events while also reducing annotation effort [43, 46]. This contrasts with traditional temporal video understanding, where actions are defined by explicit start and end times. In sports scenarios, these boundaries often become ambiguous—for example, clearly defining the precise start and end of a goal event can be challenging.
To address the ambiguity in defining temporal boundaries in sports actions, the SoccerNet dataset introduced AS as a new task specifically tailored to the sports domain [14], as illustrated in Figure 4. Alongside the dataset, S. Giancola et al. also proposed the first baseline method for AS, which involved evaluating various temporal pooling strategies on pre-extracted features from video chunks. Building on this initial work, they later introduced NetVLAD++ [34], a temporally-aware pooling module inspired by NetVLAD [47]. While standard NetVLAD aggregates frame-level features without considering temporal order—treating frames as an unordered set—NetVLAD++ explicitly incorporates temporal structure. This is particularly important in AS scenarios, where the sequence of frames before and after an event carries significant contextual information. By modeling both past and future contexts, NetVLAD++ achieved improved performance in AS.
A summary of recent methods developed for AS and PES, including their key techniques, modalities, and evaluation results on the SoccerNet benchmark, is provided in Table 2.
O. Rongved et al. [30] trained a 3D ResNet model, inspired by the architecture proposed in [48], and applied it directly to raw video frames using a 5-second sliding window approach. While 3D CNNs are effective at capturing spatiotemporal features, they are also known for their high computational cost, which can hinder real-time deployment and scalability. To address this limitation, Darwish et al. introduced the Spatio-Temporal Encoder (STE) [38], which relies solely on 1D convolutions and fully connected layers. This design significantly reduces computational complexity while maintaining strong performance, demonstrating that efficient architectures can still effectively model spatiotemporal dynamics without the overhead of 3D convolutions.
K. Vats et al. [31] proposed a multi-tower CNN architecture designed to process video information at multiple temporal scales, addressing the uncertainty in AS. In parallel, Cioppa et al. [32] introduced a context-aware loss function that explicitly models the temporal context surrounding actions. Their method takes a different approach by predicting multiple action spots within each video chunk, achieved by regressing several temporal offsets per chunk to capture the potential positions of events.
A significant performance improvement was achieved by Zhou et al. [36], who demonstrated that task-specific feature representations tailored to soccer broadcast videos can substantially enhance the effectiveness of AS models. They fine-tuned several pre-trained action recognition networks on soccer-specific data to extract high-level semantic features that better reflect the visual patterns and context of the sport. These features were then fed into Transformer-based encoding layers serving as the prediction head. Their approach achieved SOTA at the time, highlighting the importance of leveraging domain-adapted features for improving spotting accuracy in soccer video analysis.
Building on detection principles and self-supervised learning, Tomei et al. [35] presented RMS-Net, a model that combines regression and masking strategies. Inspired by object detection pipelines and pretraining techniques, RMS-Net enables the network to learn robust representations for spotting actions with improved precision.
J. Hong et al. later extended AS to PES by setting a stricter tolerance window [15]. Their proposed E2E-Spot model begins by extracting local spatial-temporal features using RegNet-Y [49] enhanced with Gate Shift Modules [50]. These features are passed through a single-layer bidirectional GRU [51], which outputs per-frame event occurrence probabilities.
In parallel, methods originally developed for TAL have shown strong performance when adapted to AS tasks. For example, J. Soares et al. proposed a detection approach based on short temporal anchors (e.g., 0.5–1.0s) where each anchor is a pair formed by a time instant and action class, predicting confidence scores and fine-grained temporal displacements for each anchor [40]. Their framework, inspired by object detection, employs both a 1D U-Net [52] and a Transformer-based encoder, with the Transformer outperforming by effectively modeling both temporal and contextual information [53].
Incorporating multi-modal input, especially audio, has also proven beneficial as audio data can contain many critical cues such as the sound when the player strikes the ball. Vanderplaetsen et al. [33] explored different fusion strategies and found that late fusion—integrating audio and visual features just before the fully connected layer—yielded the best performance.
With the rise of Transformers in computer vision [54], their application has also extended to the PES task. SpotFormer [39] adopts this approach by leveraging multiple feature extractors—including Transformer-based models (VideoMAE [55], SwinTransfrmer [56])—to extract rich spatial and temporal features from soccer videos. These features are then fed into a Transformer-based spotting network to perform precise event spotting. This multi-extractor strategy achieved strong performance, demonstrating the effectiveness of Transformer architectures in learning from diverse feature inputs. However, the use of multiple feature extractors significantly increases computational cost and inference time, making it less practical for real-time or resource-constrained applications.
To address temporal modeling more efficiently, Zhu et al. [41] proposed the use of a multi-scale Vision Transformer for feature extraction, followed by a sliding window mechanism to improve temporal understanding. These enhanced temporal features are then processed using NetVLAD++, a pooling strategy designed to aggregate local features into a robust global representation, producing the final event spotting predictions. This approach strikes a balance between temporal context modeling and computational efficiency, offering an alternative to the heavier multi-extractor pipelines.
In a similar direction, Shi et al. [37] introduced a multi-scene encoding strategy to address the variability in action durations. First, video features are extracted as a chunk and then divided into multiple subsets, each of which is processed by separate Transformer encoders. By varying chunk sizes, their model captures actions of different temporal lengths more effectively, demonstrating a flexible and adaptive approach to temporal modeling in event spotting tasks.
Building on the success of transformer and audio fusion, Xarles et al. proposed ASTRA (Action Spotting TRAnsformer) [42], a Transformer-based encoder-decoder model designed for soccer videos. ASTRA processes both visual and audio embeddings using learnable queries in the decoder, demonstrating significant performance gains on SoccerNet by leveraging multi-modal information with transformers.
Tran et al. [45] highlight a key limitation in traditional backbone models: they tend to rely heavily on global feature extraction, which often overlooks fine-grained details in the scene. As a result, these models struggle to detect events that occupy only a small spatial and temporal region of the video, such as ball positions or yellow/red card events. To address this issue, the authors proposed the Unifying Global and Local (UGL) Module, designed to capture both global environmental context and local features from task-relevant scene entities. For extracting global features, UGL adopts the same approach as the E2ESpot model [15], utilizing RegNet-Y with GSM. To capture local features, the model integrates GLIP [57], a vision-language model that processes a single image and a textual prompt describing the target objects to detect (e.g., ”ball,” ”referee,” ”card”). This represents the first use of vision-language models in PES, enabling the model to better focus on semantically relevant entities within the scene. By effectively combining global and local information, the UGL module achieved SOTA performance at the time of publication, demonstrating the efficacy of integrating vision-language models and hybrid feature modeling for fine-grained event detection in sports videos.
Recognizing the need for high temporal resolution in PES, A. Xarles et al. later introduced the Temporal Discriminability Enhancer Encoder-Decoder (T-DEED) [43]. T-DEED addresses the challenge of distinguishing between visually similar frames that occur in close succession by preserving fine-grained temporal resolution across multiple scales. It incorporates Scalable-Granularity Perception (SGP) layers [58] to boost intra-sequence feature discriminability.
Despite these advances, one major challenge persists: the reliance on large volumes of annotated data. To address this, S. Giancola et al. proposed the first active learning framework for AS [59]. Their pipeline begins with a model trained on a small labeled set, which is then used to generate predictions on unlabeled data. These predictions are ranked using uncertainty and entropy measures to identify the most informative samples. Selected clips are manually annotated and added back into the training loop. This iterative process significantly reduces the required labeled data while maintaining performance—achieving results comparable to state-of-the-art methods using only one-third of the dataset.
Most recently, J. Denize et al. proposed COMEDIAN, the first AS-specific framework that combines self-supervised learning (SSL) with knowledge distillation (KD) for initializing spatiotemporal transformers [44]. COMEDIAN features two Transformers: a spatial transformer that learns short-range context within video segments, and a temporal transformer that integrates longer-range context. The spatial transformer is trained using Momentum Contrast (MoCo) [60], a contrastive learning method that promotes feature similarity across related clips. Both components are further enhanced using a Soft Contrastive loss (SCE) [61], which distills knowledge from a bank of rich spatiotemporal features generated by a pre-trained video model. After pretraining, the entire framework is fine-tuned in a supervised manner for AS, achieving state-of-the-art results on the SoccerNet-v2 benchmark.
3.3 Evaluation Metrics
In sports video event detection, multiple evaluation metrics are employed depending on the specific task—TAL, AS, or PES. These metrics capture different aspects of detection performance, including localization precision, recall quality, and robustness across varying thresholds.
3.3.1 Mean Average Precision (mAP)
For TAL and AS tasks, the most widely used metric is Mean Average Precision (mAP), which measures how accurately models localize and classify actions. Given a ranked list of predicted intervals with associated confidence scores, precision and recall are defined as:
$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{\text{Total GT}},$ (1)
where $TP$ denotes true positives, $FP$ denotes false positives, and Total GT represents the number of ground-truth instances.
A prediction is considered a true positive if its Temporal Intersection over Union (T-IoU) with the ground truth exceeds a predefined threshold:
[…]
where $I_{p}$ and $I_{g}$ represent the predicted and ground-truth intervals, respectively.
The Average Precision (AP) for each class is calculated as:
$\mathrm{AP}=\sum_{k}(R_{k}-R_{k-1})P_{k},$ (3)
where $P_{k}$ and $R_{k}$ denote the precision and recall at the $k^{\text{th}}$ threshold. The mAP is obtained by averaging AP scores across all $C$ classes:
$\mathrm{mAP}=\frac{1}{C}\sum_{c=1}^{C}\mathrm{AP}_{c}.$ (4)
mAP offers a comprehensive evaluation by jointly considering localization and classification accuracy. It is widely accepted across different datasets, providing a common basis for benchmarking. However, it is sensitive to the choice of T-IoU thresholds (e.g., mAP@0.5 or mAP@0.75), and can penalize minor temporal misalignments heavily, which might be overly strict for fast or ambiguous actions, such as passes in soccer. Additionally, mAP alone does not separately assess the quality of action proposals independent of classification.
3.3.2 Average Recall at Average Number of Proposals (AR@AN)
In the context of TAPG, the Average Recall at Average Number of Proposals (AR@AN) is commonly employed. This metric evaluates the percentage of ground-truth instances that are successfully covered by at least one proposal, given a fixed number of proposals per video. Formally, recall is defined as:
Recall=Number of matched ground-truth instancesTotal number of ground-truth $\text{Recall}=\frac{\text{Number of matched ground-truth instances}}{\text{Total number of ground-truth instances}},$ (5)
where a match is determined based on a predefined T-IoU threshold.
AR@AN is typically reported at multiple proposal counts, such as AR@50, AR@100, or AR@200. It focuses on the completeness of the generated proposals, irrespective of the classification stage. However, it does not penalize redundancy or imprecise proposal boundaries, and therefore may overestimate performance in models that produce large or overlapping proposals.
3.3.3 Area Under the AR-AN Curve (AUC)
To assess proposal quality across a range of proposal counts, the Area Under the Curve (AUC) for the Average Recall versus Average Number of Proposals plot is calculated. Mathematically, it is defined as:
[…]
where $\text{AR}(n)$ denotes the average recall at $n$ proposals and $N$ is the maximum number of proposals considered.
AUC provides a holistic measure that captures how well a model maintains high recall across varying levels of proposal density. Models with higher AUC are generally more efficient, achieving strong recall with fewer proposals. Nevertheless, AUC can obscure detailed behaviors at specific proposal counts, which may be important in practical applications where the number of allowed proposals is constrained.
3.3.4 Tolerance Windows and AP Calculation in AS and PES
In contrast to TAL, which evaluates temporal intervals, both AS and PES treat detections as successful if the predicted timestamp falls within a predefined tolerance window centered around the ground-truth event. In AS, this window is typically broad, often ranging from 5 to 60 seconds, reflecting the nature of rapid, ambiguous actions where precise boundary annotation is difficult. PES, on the other hand, imposes a much stricter tolerance, typically constrained to 0–2 frames, requiring frame-level temporal precision.
Given this setup, the evaluation of AS and PES differs slightly from classical TAL evaluation. Rather than computing mAP over temporal intervals, AS and PES typically compute AP based on spot predictions relative to event anchors. Specifically, detections are first sorted by descending confidence scores. A prediction is considered a true positive if it falls within the defined tolerance window of any unmatched ground-truth event. Once a ground-truth event has been matched, it cannot be matched again.
For each true positive, the precision is computed at the time the true positive is found:
$\text{Precision}_{n}=\frac{\text{Number of True Positives up to }n}{\text{Number of Predictions up to }n}.$ (7)
The final AP is calculated as the average of the precision values at each true positive detection:
AP=1Total Number of Ground Truth $\mathrm{AP}=\frac{1}{\text{Total Number of Ground Truth Events}}\sum_{i=1}^{N}\text{Precision}_{i},$ (8)
where $N$ is the number of true positive detections.
This AP computation style ensures that early, high-confidence correct predictions contribute more to the final AP score, while late or incorrect predictions penalize the metric through lower precision values.
In AS, where larger tolerance windows are allowed, this formulation reflects the need for robust, high-recall detection under temporal ambiguity. However, the leniency of the large window means that models can achieve high AP even if the predicted timestamp is not highly precise. In contrast, PES evaluation, with its tight tolerance windows, demands much greater temporal accuracy; even small errors in frame prediction can result in a miss, making the task significantly harder and the AP score much more sensitive to prediction precision.
Thus, while the AP formulation for AS and PES follows a similar procedure, the strictness of the tolerance window fundamentally shifts the difficulty and interpretation of the resulting scores.
3.4 Datasets
Datasets play a critical role in supervised deep learning, providing the foundation for both model training and evaluation. Transformer-based architectures [54, 56, 62] are particularly data-dependent, often requiring large-scale, high-quality datasets to achieve strong generalization. However, annotating sports videos remains a time-intensive and expertise-driven process. For example, distinguishing between different serve types in table tennis or tennis can be highly challenging due to subtle motion variations and the high speed of play [6]. Consequently, high-quality, precisely annotated datasets are especially valuable, as sports actions often exhibit limited generalization across different contexts.
In this section, we review publicly available sports-related datasets commonly used for event detection. It is also important to note that although some datasets are originally annotated with temporal intervals, they can often be adapted for AS and PES tasks by treating the start or end frame of each interval as a discrete event label.
Table 3: Overview of sports-related datasets for event detection. Spotting denotes precise frame-level annotations, while Interval annotations specify action start and end times.
Dataset Sport Year Size / Duration Annotation Type Categories / Events
SoccerNet [14] Soccer 2018 500 videos / 764 hrs Spotting 3 (goals, cards, substitutions)
SSET [63] Soccer 2020 350 videos / 282 hrs Interval 11 event types, 15 story types
SoccerDB [64] Soccer 2020 346 videos / 669 hrs Interval 10
SoccerNet-v2 [29] Soccer 2021 500 videos / 764 hrs Spotting 17 event classes
SoccerNet Ball AS [53] Soccer 2023 7 videos Spotting 12 ball-action events
Tenniset [65] Tennis 2017 5 videos Interval 6
Tennis [15] Tennis 2022 3,345 clips Spotting 6 (court-specific ball contacts)
OpenTTGames [66] Table Tennis 2020 12 videos Spotting 3
P²A [67] Table Tennis 2024 2,721 videos / 272 hrs Interval 14 fine-grained / 8 high-level stroke classes
NCAA [68] Basketball 2016 257 videos /  1.5 hrs each Interval 14 (e.g., 3-point, dunk, steal)
Badminton Olympic [69] Badminton 2018 27 videos Interval 12
Figure Skating [17] Figure Skating 2021 11 videos Interval 4 transitions
FineDiving [70] Diving 2022 300 videos Interval 52 key pose transitions
FineGym [71] Gymnastics 2020 5,374 videos Spotting 32 fine-grained actions
MCFS [72] Figure Skating 2021 11,656 segments / 17.3 hrs Interval 130 across 4 event sets
3.4.1 Soccer
SoccerNet [14] is a large-scale dataset for soccer action localization, comprising 500 full-match videos from major European championships between 2015 and 2017. The dataset spans 764 hours of footage and occupies nearly 4TB of storage. Most event annotations are obtained by parsing official match reports. SoccerNet supports both AS and PES tasks, with manual one-second resolution annotations provided for three event types. For example, the ”card” event marks the moment a referee issues a yellow or red card due to a foul or misconduct.
SSET [63] is a smaller dataset relative to SoccerNet, containing 350 videos covering multiple soccer games, totaling 282 hours. It defines 11 event types and 15 story types. Designed primarily for TAL, its event annotations are interval-based. For instance, a ”kick” event is annotated from the moment a key player prepares to kick until the ball lands or exits the field. With slight annotation modifications, SSET can also be adapted for AS and PES.
SoccerDB [64] contains 346 high-quality soccer match videos, incorporating 270 matches from SoccerNet and 76 matches from the Chinese Super League (2017–2018) and FIFA World Cup editions. The dataset occupies 1.4TB and has a total duration of 668.6 hours. It defines 10 soccer event types with clear temporal boundaries, making it highly suitable for event detection tasks.
SoccerNet-v2 [29] builds upon SoccerNet by expanding the number of action classes from 3 to 17. Newly annotated events include ”Ball out of play,” ”Throw-in,” ”Foul,” ”Indirect free kick,” ”Clearance,” and ”Shot on target,” among others. Additionally, each timestamp is assigned a visibility tag, indicating whether the action is explicitly shown in the broadcast or inferred from gameplay context. For instance, a shot off target that is not directly shown but inferred based on subsequent play dynamics presents a unique challenge for automated detection.
SoccerNet Ball Action Spotting [53] introduces a new task focused on fine-grained ball interactions. Unlike the original SoccerNet events, which mainly targeted major game-changing actions, this task requires precise localization of frequent ball-related events, offering finer temporal granularity. The initial release included two key action classes: ”pass” (the moment the ball leaves a player’s foot) and ”drive” (the moment it is received by another player). The 2023 release provided annotations for 7 full matches, totaling 11,041 labeled timestamps. In 2024, the dataset was further expanded to include 12 distinct ball action classes, reflecting growing interest in modeling the flow of gameplay through continuous ball interactions.
3.4.2 Tennis
Tenniset [65] consists of five full-match videos from the 2012 London Olympic Games, sourced from YouTube. It defines six event types, such as ”set,” ”hit,” and ”serve,” annotated with precise temporal intervals. In addition, Tenniset provides textual descriptions of actions, such as ”quick serve is an ace,” enabling multimodal learning that combines video and text modalities.
The Tennis dataset [15], built upon the Vid2Player dataset [16], comprises 3,345 video clips from 28 professional tennis matches recorded at 25 or 30 FPS. Events are categorized into six classes, including ”player serve ball contact,” ”regular swing ball contact,” and ”ball bounce,” further divided based on court side (near or far court). The dataset supports fine-grained action spotting in tennis and facilitates evaluation under strict temporal precision settings.
3.4.3 Table Tennis
OpenTTGames [66] consists of 12 high-definition table tennis match videos recorded at 120 FPS, containing 4,271 labeled events. The dataset includes three event types: ball bounces, net hits, and empty events, all annotated with frame-level precision. OpenTTGames is particularly suited for training models targeting bounce event detection under high-speed gameplay conditions.
P2A [67] is one of the largest publicly available table tennis datasets, comprising 2,721 untrimmed broadcast videos with a total length of 272 hours. The dataset features footage from major international and Chinese tournaments, including the Tokyo 2020 Summer Olympics, World Cups, and World Championships. Annotations were conducted in collaboration with professional players and referees, defining 14 fine-grained stroke classes, which are further grouped into 8 higher-level action categories. Each action is annotated with precise start and end frames, enabling accurate event localization. The reliability of the dataset is strengthened by professional validation of the annotations.
3.4.4 Other Sports
NCAA [68] consists of 257 untrimmed college basketball game videos, each approximately 1.5 hours long. The dataset provides 14,548 video segments, with precise start and end times for 14 action categories, supporting temporal action localization tasks.
Badminton Olympic [69] contains 27 badminton match videos sourced from the official Olympic YouTube channel. It includes time-interval annotations for 12 action types, such as ”serve,” ”backhand,” and ”smash,” as well as point-level annotations, making it suitable for both action spotting and match-level analysis.
BadmintonTrack [73] is another badminton dataset, comprising 77,000 annotated frames from 26 unique singles matches filmed from an overhead broadcast-view camera. Originally, timestamp information indicating when a player struck the shuttlecock was included [74], although this metadata has since been removed in updated versions.
The Figure Skating dataset [17] contains 11 videos recorded at 25 FPS, featuring 371 short program performances from the 2010–2018 Winter Olympics and the 2017–2019 World Championships. It defines four transition event types critical for evaluating temporal precision in figure skating analysis.
FineDiving [70] includes 300 professional diving videos sourced from major competitions such as the Olympics, World Cup, and World Championships. The dataset defines 52 fine-grained action types, along with 29 sub-action types and 23 difficulty levels, offering a comprehensive benchmark for procedural action quality assessment.
FineGym [71] provides 5,374 gymnastics performances from international competitions. Each video is annotated with a hierarchical structure categorizing 32 spotting classes, covering disciplines such as balance beam and floor exercise, enabling fine-grained action spotting and classification.
MCFS [72] is a large-scale figure skating dataset comprising 11,656 video segments across 38 competitions, totaling 17.3 hours and 1.7 million frames. Annotations follow a hierarchical structure of 4 high-level action sets, 22 subsets, and 130 element actions, making MCFS highly suitable for dense temporal action localization tasks in figure skating.
4 Challenges and Future Directions
In this section, we critically examine current challenges in sports event detection and outline specific, actionable future research directions to address these limitations.
4.1 Generalization Across Diverse Sports
Most existing AS and PES models are developed and validated primarily within individual sports contexts, such as soccer (e.g., SoccerNet). These specialized approaches perform well within their target sports but often exhibit limited capability for generalization, as they rely on domain-specific characteristics like camera angles, event definitions, and player interactions.
Only a limited number of approaches have systematically evaluated their effectiveness across multiple sports datasets (see Table 2), including tennis, figure skating, diving, and gymnastics. This gap highlights a significant limitation: current models struggle to generalize to sports with different visual characteristics, dynamics, and temporal structures. Future research must prioritize:
  • Development of universal AS and PES models leveraging cross-domain transfer learning, possibly employing large-scale pre-trained vision-language models (e.g., CLIP [75]), to enhance cross-sport adaptability.
  • Rigorous benchmarking across diverse sports to validate generalizability, robustness, and practical utility in varied real-world scenarios.
  • The creation and release of new annotated datasets spanning a wider range of sports—particularly underrepresented domains such as volleyball, hockey, fencing, and athletics—to support broader generalization and fair evaluation.
4.2 Advancements in Unsupervised and Low-Supervision Methods
Creating extensive labeled datasets for sports event detection is costly, labor-intensive, and often demands expert domain knowledge, particularly in technical sports such as gymnastics, tennis, and figure skating. To mitigate these barriers, recent methods have begun exploring low-supervision learning paradigms such as knowledge distillation and active learning [44, 59]. These approaches significantly reduce dependency on large annotated datasets by transferring knowledge from pretrained models or selectively labeling informative samples.
Nevertheless, fully unsupervised or self-supervised methods remain largely unexplored. Future research directions include:
  • Development of self-supervised frameworks that leverage temporal consistency, contrastive learning, and multimodal alignment to extract meaningful event representations from unlabeled or weakly labeled video data.
  • Integration of unsupervised learning with domain adaptation techniques to further boost model performance and generalization across multiple sports domains.
These strategies will be essential for building scalable, efficient, and broadly applicable event detection models suitable for diverse sports analytics.
4.3 Enhanced Multimodal Fusion Approaches
Although existing AS and PES methodologies predominantly rely on visual data, audio cues can significantly enrich the detection of critical events in sports. Examples include sounds of ball impacts, crowd reactions, or figure skaters landing on ice—all of which provide valuable complementary temporal signals.
Currently, multimodal models predominantly employ straightforward fusion methods such as concatenation or late fusion [33, 42], which do not fully capture the complex interactions between modalities. To address these limitations, future research should explore more sophisticated fusion techniques:
  • Attention-based cross-modal transformers and gated attention mechanisms, capable of dynamically weighting and integrating audio and visual modalities.
  • Modality-specific encoders coupled with temporal alignment mechanisms to capture precise event timings, especially in noisy or visually ambiguous scenarios.
  • Exploiting commentary audio through automatic speech recognition and natural language processing to provide additional semantic context and weak supervision signals, aligning spoken descriptions directly with visual event occurrences.
Advancing multimodal fusion methodologies represents a crucial opportunity to significantly enhance the accuracy, robustness, and practical applicability of AS and PES frameworks in sports video analytics.
5 Conclusion
In this survey, we presented a comprehensive review of the current state of deep learning-based video event detection, with a particular focus on the critical tasks of TAL, AS, and PES within the sports analytics domain. By systematically categorizing and evaluating recent datasets and advanced methodologies, we identified key limitations, including poor generalizability across diverse sports contexts, reliance on extensive annotated datasets, and the underutilization of multimodal information sources.
Our analysis highlights the urgent need for the development of generalized frameworks capable of robust performance across multiple sports disciplines, potentially leveraging transfer learning and cross-domain evaluation strategies. Moreover, advancing unsupervised and low-supervision approaches, such as self-supervised learning and active learning, will be critical to reducing annotation costs and enabling scalable deployment. In parallel, future research should explore sophisticated multimodal fusion techniques that effectively integrate visual, audio, and textual modalities to enhance temporal precision, robustness, and event understanding in complex real-world environments.
By addressing these challenges, future efforts can significantly advance the capabilities of automated sports video analytics, leading to more comprehensive, accurate, and efficient systems capable of delivering actionable insights across a wide range of sports applications.