### v2

# Action Spotting and Precise Event Detection in Sports: Datasets, Methods, and Challenges

Abstract
Video event detection is central to modern sports analytics, enabling automated understanding of key moments for performance evaluation, content creation, and tactical feedback. While deep learning has significantly advanced tasks like Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES), existing surveys often overlook the fine-grained temporal demands and domain-specific challenges posed by sports. This survey first provides a clear conceptual distinction between TAL, AS, and PES, then introduces a methods-based taxonomy covering recent deep learning approaches for AS and PES, including feature-based pipelines, end-to-end architectures, and multimodal strategies. We further review benchmark datasets and evaluation protocols, identifying critical limitations such as reliance on broadcast-quality footage and lenient multi-label metrics that hinder real-world deployment. Finally, we outline open challenges and future directions toward more temporally precise, generalizable, and practical event spotting in sports video analysis.
1 Introduction
Sports represent one of the largest global markets, projected to reach 599.9 billion US dollars by 2025 and 826 billion US dollars by 2030, with a compound annual growth rate of 6.6% [1]. Beyond its economic impact in industries such as media, marketing, and apparel, sports are fundamentally focused on athletic performance, where optimising player efficiency, refining game strategies, and enhancing fan engagement are critical. The rise of sports analytics, the systematic collection, processing, and analysis of performance data, has enabled data-driven decision-making, leading to fundamental changes in strategy. For instance, basketball has seen an increased reliance on three-point shooting, guided by predictive models that estimate expected points from various court locations [2].
Within this context, video event detection has emerged as a fundamental yet challenging task in sports analytics. Accurate identification of key moments, such as corner kicks in soccer, rally conclusions in racket sports, or scoring events across disciplines, provides critical insight to coaches and athletes, supporting more effective performance analysis and tactical decision-making. Moreover, event detection benefits downstream applications by filtering non-playing segments, optimising computational resources for subsequent tasks such as object tracking, and enabling automated highlight generation for commercial broadcasting and fan engagement.
Computer Vision (CV) has played a pivotal role in advancing sports video analysis, enabling automated player tracking, action recognition, tactical analysis, injury prevention and officiating through advanced visual data processing [3, 4, 5, 6, 7, 8]. While traditional CV methods laid the foundation for many applications [5, 9], they heavily relied on handcrafted features and struggled with dynamic environments, occlusions, and real-time constraints. Recent breakthroughs in Deep Learning (DL), particularly through Convolutional Neural Networks (CNNs) and Transformer architectures, have significantly enhanced the accuracy, efficiency, and scalability of sports video analysis tasks [5]. These modern techniques now power real-time object detection, robust player tracking such as [10, 11, 12], pose estimation [13], and fine-grained event recognition [14, 15], fundamentally reshaping the landscape of sports analytics.
Figure 1: Annual publication count from 2010 to 2024 based on a Scopus keyword search for ”sports” AND ”deep learning”.
To further illustrate the growing interest in this domain, we compiled a publication trend analysis using Scopus, based on keyword searches related to ”sports” and ”deep learning.” The resulting bar chart (Figure 1) shows a clear year-on-year increase in the number of relevant publications from 2010 to 2024, reinforcing the expanding research momentum in this field.
As video event detection in sports continues to evolve with the rise of DL-based CV techniques, a systematic literature review is essential to consolidate recent progress, clarify emerging task definitions—such as Temporal Action Localization (TAL), Action Spotting (AS), and Precise Event Spotting (PES)—and identify current methodological gaps and research opportunities.
Table 1: Comparison of Temporal Action Localisation (TAL), Action Spotting (AS), and Precise Event Spotting (PES).
Aspect TAL AS PES
Output Type Temporal interval Single key frame Single key frame
Annotation Format Start and end times Single timestamp Single timestamp
Tolerance Window $\sim$1–5 seconds 5–60 frames 0–2 frames
Best Suited For Long-duration actions Ambiguous, fast-paced actions Frame-accurate event detection
Annotation Cost High Medium High
Use Cases Generic action localization Sports highlight detection Fine-grained critical events
Previous studies, such as those by Ghosh et al. [5], broadly survey AI applications in sports analytics but do not specifically address DL-based CV methods. Similarly, Thomas et al. [9] focus primarily on traditional CV approaches designed for multi-camera systems. In contrast, our survey specifically targets deep learning-based approaches—including recent advances in CNNs and Vision Transformers—for event detection tasks within monocular video contexts, enhancing its relevance for real-world deployment. Other comprehensive surveys by Naik et al. [3], Zhao et al. [4], Kamble et al. [16], and Wu et al. [6] extensively review CV methods across various sports analytics tasks, including object tracking and action recognition. However, they do not specifically address video event detection. The most closely related work by Hu et al. [17] provides an extensive overview of TAL methods, but it does not cover the increasingly critical tasks of AS and PES, nor is it specifically focused on the sports domain.
This survey addresses these significant gaps by making the following key contributions:
  • Task-Oriented Taxonomy and Analysis: We introduce a structured taxonomy that formally defines and distinguishes AS and PES, positioning them within the broader landscape of video event detection. This taxonomy emphasizes the unique temporal and semantic challenges of sports analytics.
  • In-Depth Methodological Survey: We comprehensively review deep learning-based approaches for AS and PES, spanning feature-based pipelines, end-to-end architectures, and multi-modal frameworks leveraging convolutional, recurrent, and transformer-based designs.
  • Evaluation Protocols and Dataset Landscape: We summarize key benchmark datasets and evaluation metrics used across TAL, AS, and PES tasks, and critically examine their suitability for real-world deployment, including limitations in confidence thresholding and multi-label predictions.
  • Critical Insights and Future Directions: We identify open challenges such as poor generalization across diverse sports, limited supervision strategies, and unrealistic evaluation schemes. We propose actionable directions to guide the development of more robust and deployable spotting models.
By clearly delineating these specific and timely contributions, our survey offers a distinct and valuable perspective within the broader sports analytics literature.
The remainder of the paper is organized as follows. Section 2 provides clear definitions of the various tasks involved in sports video event detection. In Section 3, we review existing methodologies developed for this domain. Section 4 and 5 introduce relevant evaluation metrics and datasets commonly used for sports event detection. Section 6 outlines key challenges and future research directions. Finally, Section 7 concludes the paper with a summary of insights and findings.
Figure 2: Example of Temporal Action Localization: in tennis, the time interval of the whole serve action has been labelled (blue bar).
Figure 3: Example of Precise Event Spotting: in table tennis, the moment a player contacts the ball during a serve (red) or when a ball bounces on the table (blue).
2 Sports Event Detection
In sports video analysis, three key tasks have emerged for temporal event detection: Temporal Action Localisation (TAL), AS, and PES. While related, they differ in output format, annotation granularity, and application focus (see Table 1).
2.1 Temporal Action Localisation
TAL—also referred to as Temporal Action Detection (TAD)—aims to detect and classify action segments within untrimmed videos. It builds on Temporal Action Proposal Generation (TAPG), which identifies candidate segments likely to contain actions, without assigning labels [18, 19, 17].
Although originally designed for general activity understanding, many TAL methods have been adapted to sports due to their strength in modeling temporal structure. TAPG is especially challenging because it requires precise boundary prediction, while classification within those segments is typically more accurate once good proposals are established [17].
2.2 Action Spotting
AS, introduced by Giancola et al. [20], addresses the difficulty of defining precise action boundaries in sports, where events are fast, overlapping, or continuous. Instead of annotating start and end times, AS uses a single timestamp to represent an event, simplifying annotation and avoiding ambiguity.
While both TAL and AS seek to locate events temporally, AS trades interval-level precision for a more efficient and scalable annotation strategy, better suited for rapid and ambiguous sports actions.
2.3 Precise Event Spotting
PES extends AS by enforcing strict frame-level precision. In sports like tennis [21] and figure skating [22], even a 1–2 frame error can miss critical events. Similarly, feedback from Table Tennis Australia confirms that pinpoint accuracy is essential for detecting ball contact or bounce points, where the ball may travel across frames in under 100 milliseconds.
PES thus requires event detection within tight tolerance windows (typically 1–5 frames), making it ideal for high-precision, frame-accurate tasks in elite sports contexts.
3 Video Event Detection
In this section, we provide a comprehensive overview of video event detection methods, covering both general-purpose approaches and those specifically designed for sports. We begin by reviewing foundational techniques originally developed for generic activity detection, many of which have influenced early progress in sports video analysis. We then shift our focus to sports-specific methodologies, including the recent emergence of tasks such as AS and PES, which address the unique challenges of detecting fine-grained events in fast-paced, ambiguous sports scenarios.
3.1 General Video Event Detection
Current methods for TAL typically fall into two main categories: Global-to-Local and Local-to-Global approaches. Global-to-Local methods generate action proposals through sliding windows or pre-defined anchor structures, segmenting videos into fixed-length or hierarchically structured intervals to localise actions. Local-to-Global methods independently evaluate each frame to predict the likelihood of action start, action end, and overall actionness. Proposals are subsequently created by connecting high-confidence start and end frames, followed by a confidence-based evaluation to confirm whether the proposal contains an action. The difference between these two approaches is shown in Figure 4. The subsequent sections will explore these two approaches comprehensively, discussing their strengths, limitations, and practical applications in sports event detection.
Figure 4: Comparison of Global-to-Local (GTL, left) and Local-to-Global (LTG, right) approaches. In GTL, predefined temporal anchors are classified as action or background, followed by a regressor that refines the event’s time interval. In contrast, LTG first predicts per-frame start and end probabilities to define the temporal boundaries and then classifies the resulting interval.
3.1.1 Global to Local
J. Gao et al. proposed the Temporal Unit Regression Network (TURN) to address the temporal action proposal generation problem [23]. TURN first decomposes an untrimmed video into short video units, which serve as fundamental processing blocks, with visual features extracted from each unit using a base model C3D [24]. To construct meaningful clip-level representations, TURN pools a set of contiguous units across multiple temporal scales, forming a clip pyramid. Each clip is further enriched by concatenating features from both its internal and surrounding units, ensuring contextual information is preserved. At this stage, each clip is treated as a proposal candidate, and TURN outputs a confidence score indicating whether the clip contains an action instance, along with two regression offsets that refine the predicted start and end times of the action within the clip. Finally, Non-Maximum Suppression (NMS) is applied to eliminate redundant proposals, improving localization accuracy and ensuring high-quality action boundaries. They have also proposed a new metric for this task which is called Average Recall vs. Frequency of retrieved proposals (AR-F) which covers two criteria, (1) it should be capable of evaluating the performance of different methods on the same dataset efficiently; (2) it should be capable of evaluating the performance of the same method across different dataset. Even though TURN has achieved the SOTA performance on two challenging datasets THUMOS-14 [25] and ActivityNet [26], with the highest AR at the time because of the sliding window nature where it covers all segments in the video however the temporal boundaries are not precise enough, indicating its flaws when accurately deciding the correct time frame for the event [27, 19, 28].
On the other hand, Temporal Action Grouping (TAG) [29] was proposed as an alternative to fixed temporal scale methods, drawing inspiration from the watershed algorithm [30]. TAG groups continuous high-score regions into proposals rather than relying on predefined temporal scales. It applies a binary classification to each unit or snippet (a few contiguous frames) to generate actionness scores, which indicate the likelihood of an action occurring in that unit. This method is capable of producing high-precision temporal boundaries, provided that the actionness scores are accurate. However, its performance heavily depends on the quality of the predicted actionness scores, making it less effective when predictions are unreliable [27].
To address these limitations, the Complementary Temporal Action Proposal (CTAP) was introduced by J. Gao et al. [27], leveraging the strengths of both actionness-based proposals and sliding-window proposals. CTAP first generates both types of proposals and then applies a proposal complementary filter, which collects missing correct proposals that TAG alone might overlook by retrieving them from the sliding-window proposals. In this stage, a binary classifier is used, taking a sequence of unit features as input and outputting a probability score that indicates whether a proposal can be accurately detected by the unit-level actionness model and TAG. Finally, a Temporal Convolutional Adjustment and Ranking Network is employed to rank proposals and refine their boundaries. Experimental results on THUMOS-14 and ActivityNet v1.3 demonstrate that CTAP outperforms TURN in AR, highlighting its effectiveness in improving proposal quality. However, the combination of multiple proposal generation methods can impact computational efficiency, potentially increasing processing time.
GTL methods follow a two-stage paradigm similar to R-CNN: they first generate candidate proposals and then classify them. This approach has shown strong performance in early temporal action localization work. However, a key limitation lies in the fixed-length nature of the proposals, which does not align well with the highly variable durations of action instances across different videos. As a result, these methods often rely on dense sliding windows to ensure sufficient coverage, leading to high computational overhead. Furthermore, while classification accuracy is generally strong, GTL approaches often struggle to produce precise temporal boundaries, limiting their effectiveness for fine-grained event localization [17].
Figure 5: Examples of events defined in the context of soccer. From top to bottom: Card—the referee shows a card to a player; Substitution—the moment a player enters the field; and Goal—the ball crosses the goal line. All examples are from SoccerNet [20].
3.1.2 Local to Global
T. Lin et al. proposed the Boundary Sensitive Network (BSN), the first approach designed to predict the start, end, and actionness of each frame in a video [18]. Proposals are generated by pairing start and end points with high probabilities, while low-confidence candidates are discarded by an evaluation module. BSN significantly outperforms previous methods and achieves high performance even with relatively few proposals. On the THUMOS14 dataset, BSN surpasses all SOTA methods across a range of proposal counts (10 to 1000). Notably, when the average number of proposals is 50, BSN improves AR from 21.86% to 37.46%, demonstrating its effectiveness in capturing action boundaries with fewer proposals.
However, BSN suffers from inefficiencies due to its proposal feature construction and confidence evaluation processes, which are applied individually to each proposal, leading to high computational overhead. To address this issue, Boundary-Matching Network (BMN) was introduced [19], incorporating a Boundary-Matching (BM) mechanism for more efficient confidence evaluation of densely distributed proposals. Instead of evaluating each proposal separately, BMN represents proposals as matching pairs of their start and end boundaries and constructs a two-dimensional BM confidence map. This map models densely distributed proposals by encoding continuous starting boundaries and temporal durations, allowing confidence scores to be computed simultaneously for all proposals. By leveraging this structure, BMN significantly improves efficiency while maintaining strong performance in temporal action proposal generation.
Building upon these advancements, BSN++ was introduced as an enhanced version of BSN to further refine proposal generation and confidence evaluation [28]. While BSN effectively captures action boundaries, it lacks a mechanism to assess the completeness of proposals, often leading to inaccurate segmentations. BSN++ addresses this by incorporating a proposal completeness modeling module, which evaluates whether a detected segment fully encapsulates an action instance. Additionally, BSN++ enhances feature representation by introducing context-aware proposal features, ensuring that both internal and surrounding contextual information contribute to more precise action localization. To improve proposal ranking, a global-local confidence fusion module is integrated, leveraging both local proposal confidence and global consistency to reduce false positives. Furthermore, BSN++ adopts the BM confidence map to evaluate multiple proposals simultaneously, significantly improving computational efficiency. These refinements enable BSN++ to generate higher-quality proposals while reducing redundancy, leading to superior performance in temporal action proposal generation.
However, Z. Qing et al. argue that existing methods struggle to effectively model temporal relationships, as neither convolution-based approaches nor global fusion mechanisms can fully capture long-range dependencies. Additionally, relying solely on internal or boundary context for regression is insufficient for generating proposals with precise boundaries [31]. To address these limitations, they proposed the Temporal Context Aggregation Network (TCANet) for high-quality proposal generation, which consists of two key components. First, the Local-Global Temporal Encoder (LGTE) is introduced to capture both local and global temporal relationships in a channel-grouping fashion. The Local Temporal Encoder (LTE) is responsible for processing the first A groups of the input feature maps, which are linearly transformed and evenly divided into N groups along the channel dimension to enhance short-term temporal modeling. Simultaneously, the remaining N - A groups are handled by the Global Temporal Encoder (GTE), which captures long-range dependencies. By combining these two encoders, LGTE effectively integrates long-term contextual information while preserving structural and fine-grained local details, leading to richer proposal representations. Second, the Temporal Boundary Regressor (TBR) is designed to refine proposal boundaries by leveraging both internal and boundary context. This is achieved through a dual regression approach: (1) Frame-level boundary regression, which refines the starting and ending locations of proposals with boundary sensitivity, and (2) Segment-level boundary regression, which adjusts the center location and duration of proposals based on a holistic perception of their structure. The final predicted proposal boundaries are obtained by fusing the outputs of both regression modules, ensuring precise action localization. Experimental results demonstrate that TCANet significantly enhances action proposal generation and action detection, outperforming existing methods by improving both boundary accuracy and proposal completeness.
Previous works typically use tIoU between the proposal and the ground truth action instance—referred to as the action score—to evaluate proposal confidence and develop high-quality temporal action proposals. However, H. Yang et al. argue that background information is also significant but was previously overlooked. Motivated by this observation, they proposed the Background Constraint Network (BCNet) to generate high-quality temporal action proposals [32]. BCNet consists of two main modules: the Action-Background Interaction (ABI) Module and the Boundary Prediction (BP) Module. Inspired by the Transformer [33], the ABI module first employs self-attention to capture relationships between frames and enhance feature representations. The enhanced frame features are then processed by the Difference-Attention Unit, which aggregates features based on the inconsistency between action and background features under supervision. Finally, a prediction head encodes both action and background features, followed by a Sigmoid layer to generate action and background probability sequences. During training, unlike previous works that rely solely on action score loss, BCNet incorporates background score loss to further refine proposal confidence. Extensive experiments on the ActivityNet-v1.3 and THUMOS14 datasets demonstrate the effectiveness of explicitly modeling background information and leveraging attention mechanisms in TAPG.
Compared to GTL, LTG methods address the limitation of fixed proposals by enabling flexible boundary generation through frame-level classification. This allows for more adaptive localization of action segments. However, LTG approaches are often sensitive to fluctuations in boundary predictions, which can result in unstable or fragmented proposals [17].
Figure 6: Illustration of three representative modeling strategies for AS and PES tasks. Pooling-Based Models perform temporal aggregation on extracted features; Encoder-Based Models use temporal encoders (e.g., 1D/3D CNNs, Transformers) to capture sequential dependencies; and Frame-Discriminative Models modify the backbone to integrate frame-level temporal reasoning directly for improved spotting accuracy.
Figure 7: Generalized architecture for AS and PES models. Most approaches follow a four-stage modular pipeline: input sampling (frame-wise or chunk-based), feature extraction (via 2D/3D CNNs or transformers), temporal modeling (e.g., 1D CNNs, RNNs, or self-attention), and classification. This structure underlies a variety of modeling strategies across the AS and PES literature.
3.2 Sports Video Event Detection
While general video event detection methods laid the foundation for sports analytics, the unique demands of sports—such as rapid actions, ambiguous event boundaries, and the need for frame-level precision—have driven the development of specialized approaches. TAL methods, originally designed for general activity understanding, often fall short in these contexts due to their reliance on coarse temporal intervals.
To address these limitations, two sports-specific tasks have emerged: AS and PES. Unlike TAL, which focuses on identifying action segments, AS and PES aim to detect the exact moment an event occurs. This shift from segment-based to keyframe-based detection better reflects the temporal granularity required in sports settings.
Although AS and PES differ in evaluation metrics and frame-level precision requirements, many detection models are applicable to both. We categorize existing approaches according to their underlying architectural strategies:
  • Pooling-Based Methods aggregate features from pretrained models within a fixed temporal window using different pooling techniques.
  • Temporal Encoder-Based Methods employ sequence models (e.g., 1D/3D CNNs or Transformers) to model temporal dependencies across frames.
  • Frame-Discriminative Models focus on designing architectures to capture fine-grained frame-level differences, enabling more precise temporal localization tailored to the demands of spotting tasks.
  • Multi-Modal Fusion Methods incorporate non-visual modalities, such as audio, to enhance detection performance.
  • Data-Efficient Learning Approaches reduce reliance on large-scale annotations through active learning, self-supervised learning, or knowledge distillation.
Figure 6 presents a visual overview of three representative architectural strategies—Pooling-Based Models, Encoder-Based Models, and Frame-Discriminative Models—that form the core of most AS and PES approaches. Multi-modal methods typically adopt one of these architectures while incorporating additional input sources such as audio or text. In contrast, data-efficient learning approaches (e.g., self-supervised or active learning) represent a separate paradigm focused on training strategy rather than model structure, and are thus not easily categorized within the same architectural taxonomy.
While the taxonomy in Figure 6 organizes AS and PES methods by high-level architectural strategy, most models—regardless of category—follow a shared modular pipeline. This typically consists of input sampling, frame-level feature extraction, temporal modeling, and classification. Figure 7 illustrates this generalized structure, which underpins both frame-wise and chunk-based approaches, and helps contextualize where architectural variations occur within the overall flow.
3.2.1 Pooling Based Approaches
Pooling-Based Approaches aggregate features over fixed temporal windows using techniques like average pooling, NetVLAD, or its temporally-aware variant NetVLAD++. These models typically use 2D or 3D CNNs to extract frame-level features, followed by simple aggregation and classification.
One of the earliest works in this category is the SoccerNet baseline [20], which used average and max pooling on pre-extracted features from C3D [24], I3D [34], and ResNet [35]. Rongved et al. [36] later extended this with a 3D ResNet—adapted from [37]—using a 5-second sliding window to better capture temporal dynamics. NetVLAD++ [38] further improved upon these ideas by explicitly modeling temporal order in the aggregation process, enabling better contextual understanding of past and future frames.
Zhou et al. [39] advanced this line of work by fine-tuning multiple action recognition backbones—including TPN [40], GTA [41], VTN [42], irCSN [43], and I3D-Slow [44]—on soccer video snippets. The combined features, when processed by NetVLAD++ or Transformer-based encoders, achieved state-of-the-art performance on the SoccerNet benchmark.
Zhu et al. [45] introduced a more efficient approach by utilizing a single multi-scale Vision Transformer for feature extraction, followed by NetVLAD++ pooling. This model strikes a balance between temporal modeling capacity and computational efficiency, making it more suitable for deployment in resource-constrained settings.
To better leverage the structure of events over time, Cioppa et al. [46] proposed a context-aware loss function that dynamically weights frames based on their proximity to annotated events. This smooth temporal weighting significantly improved focus on relevant cues, resulting in a 12.8% performance boost on SoccerNet-v1. However, its effectiveness diminishes in denser datasets like PES [47], which require higher frame-level precision.
3.2.2 Encoder-Based Methods
Encoder-Based Methods enhance feature exploitation pipelines by replacing pooling operations with sequence models that learn temporal dependencies directly from the input features. These models include 1D CNNs, 3D CNNs, and Transformers, which allow for richer contextual modeling.
To address variability in action duration, Vats et al. [48] proposed a multi-tower CNN architecture operating at multiple temporal resolutions. This multi-scale design captures both short-term and long-range dependencies, improving robustness across diverse sporting actions.
Tomei et al. [49] introduced RMS-Net, a hybrid model combining regression-based localization with a masking strategy inspired by object detection and self-supervised learning. This design promotes the learning of temporally structured signals, improving the precision of event boundaries.
SpotFormer [50] further demonstrates the strength of encoder-based models by fusing features from multiple pretrained backbones (e.g., VideoMAE [51], Swin Transformer [52]) and processing them with a Transformer spotting head. While highly accurate, this approach is computationally intensive and less suitable for real-time use.
To address computational constraints, Darwish et al. [53] proposed the Spatio-Temporal Encoder (STE), a lightweight model employing 1D convolutions and MLPs. Despite its simplicity, STE delivers competitive performance by efficiently capturing key temporal patterns.
Soares et al. [54] adapted an anchor-based detection framework—originally designed for TAL—to AS. Their model combines a 1D U-Net with a Transformer encoder, effectively capturing both local and long-range temporal dependencies, as reflected in their strong performance on SoccerNet Challenge [55].
Shi et al. [56] addressed variable-length events through a multi-scene encoding strategy, segmenting video chunks into variable-length subsets processed by separate Transformers. This enables the model to adaptively capture temporal context across varying durations.
Table 2: Summary of methods for Action Spotting (AS) and Precise Event Spotting (PES). Performance is grouped under the Test and Challenge sets. All results are reported on the SoccerNet Action Spotting dataset, where italicized entries indicate results from SoccerNet-v1 [20], and plain text entries are from SoccerNet-v2 [57]. Bold numbers indicate the highest scores in each column. A ✓in the last column means the method was evaluated on datasets beyond SoccerNet. C.D. Eval means cross-dataset evaluation.
Method Year Category Parameter Size Test Set Challenge Set C.D Eval.
Tight Loose Tight Loose
Giancola et al. [20] 2018 Pooling-Based 31.37 30.74
Rongved et al. [36] 2020 End-to-End Temporal Learning 56.3
Vats et al. [48] 2020 Temporal Encoder-Based 60.1
CALF [46] 2020 Pooling-Based 41.61 42.22
Vanderplaetsen et al. [58] 2020 Multi-Modal Fusion 39.90
NetVLAD++ [38] 2021 Pooling-Based 53.40 52.54
RMS-Net [49] 2021 End-to-End Temporal Learning 63.49 60.92
Zhou [39] 2021 Pooling-Based 47.05 74.77 49.56 74.84
E2E-Spot [47] 2022 End-to-End Temporal Learning 4.5M 66.73 73.62
Shi et al. [56] 2022 Temporal Encoder-Based 55.20
STE [53] 2022 Temporal Encoder-Based 2.3M 58.29 71.58 58.71 70.49
SpotFormer [50] 2022 Temporal Encoder-Based 60.90 81.50
Soares et al. [54] 2022 Temporal Encoder-Based 8.9M 65.07 78.59 68.33 78.06
Zhu et al. [45] 2022 Pooling-Based 52.04 60.86
ASTRA [59] 2023 Multi-Modal Fusion 70.10 79.21
T-DEED [60] 2024 End-to-End Temporal Learning 16.4M
COMEDIAN [61] 2024 Data-Efficient Learning 29.1M 73.10 68.38 73.98
Tran et al. [62] 2024 End-to-End Temporal Learning 62.49 73.98 69.38 76.15
3.2.3 Multi-Modal Based Methods
Multi-Modal Based Methods extend visual modeling by incorporating complementary modalities, such as audio. Acoustic cues—like whistles, ball strikes, or crowd reactions—often correlate with event boundaries and offer additional context.
Vanderplaetsen et al. [58] evaluated several fusion strategies and found that late fusion, where audio and visual features are combined just before classification, yielded the best performance on SoccerNet.
Expanding on this, Xarles et al. [59] introduced ASTRA, a Transformer-based encoder-decoder model that jointly processes audio and visual embeddings using learnable cross-modal queries. ASTRA achieved strong performance, showcasing the power of cross-modal attention in sports event detection.
3.2.4 Frame-Discriminative Models
Frame-Discriminative Models aim to enhance the spatiotemporal representational capacity of networks by modifying the base architecture or training strategy, often incorporating frame-specific mechanisms to improve temporal discriminability and support precise event spotting. Unlike pooling and encoder based methods, which rely on models designed for unrelated tasks, these approaches allow the network to jointly learn low-level visual features and high-level temporal semantics tailored to the spotting task, resulting in more discriminative and temporally precise predictions.
Hong et al. [47] proposed E2E-Spot, an end-to-end trainable model built on RegNet-Y for visual encoding. To capture fine-grained temporal dynamics, they integrated Gate Shift Modules (GSM) [63], which explicitly model temporal shifts at the frame level. A bidirectional GRU [64] is then employed to process the sequential features and generate per-frame event probabilities, enabling frame-accurate localization suitable for PES.
To overcome the limitations of global-only modeling in E2E-Spot, Tran et al. [62] introduced the Unifying Global and Local (UGL) module. While maintaining the RegNet-Y + GSM backbone for global spatiotemporal encoding, UGL incorporates GLIP [65], a vision-language model, for fine-grained local semantic reasoning (e.g., detecting contextual cues like “referee” or “ball”). This fusion of global temporal encoding with localized semantic awareness enhances the model’s ability to detect subtle or ambiguous events such as fouls or off-screen actions, achieving state-of-the-art results on SoccerNet.
To further improve temporal resolution and disambiguate closely spaced events, Xarles et al. proposed T-DEED [60], a Transformer-based encoder-decoder architecture specifically designed for PES. It introduces Scalable-Granularity Perception (SGP) layers [66] and Gate Shift Fusion (GSF) modules [67], enabling the model to retain temporal discriminability across tightly clustered frames. This design proves especially effective in scenarios with rapid, visually similar events—common in fast-paced sports settings requiring high precision.
3.2.5 Other Models
Despite significant progress, a key bottleneck in AS and PES remains the reliance on large-scale annotated datasets. To address this, Giancola et al. [68] proposed the first active learning framework for action spotting. Their approach begins with a model trained on a small labeled subset, which is then used to generate predictions on unlabeled data. These predictions are ranked by uncertainty and entropy-based heuristics to identify the most informative samples. Manually annotated clips are iteratively added back into the training pool, reducing annotation effort while maintaining competitive performance—achieving results comparable to state-of-the-art models with only one-third of the labeled data.
More recently, Denize et al. introduced COMEDIAN [61], the first AS-specific framework to combine self-supervised learning (SSL) and knowledge distillation (KD) for pretraining spatiotemporal Transformers. The architecture includes a spatial transformer that captures short-range local context and a temporal transformer that models long-range dynamics. The spatial component is trained using Momentum Contrast (MoCo) [69], encouraging consistent representations across temporally adjacent clips. Additionally, a Soft Contrastive Loss (SCE) [70] is used to distill knowledge from a rich feature bank generated by a pre-trained video model. After the SSL phase, the network is fine-tuned with supervision on the AS task, yielding state-of-the-art performance on SoccerNet-v2 while reducing reliance on labeled data.
While existing methods span a range of architectural strategies, each has trade-offs. Pooling-based methods offer computational simplicity by aggregating features over time, though this comes at the cost of temporal precision. Temporal encoders improve contextual modeling but often rely on generic, non-specialized features. Multi-modal methods enhance detection with audio or text cues but depend on modality alignment, while data-efficient approaches reduce labeling costs yet remain limited by their supervision strategies.
A common limitation is the reliance on backbones designed for unrelated tasks like image classification or coarse action recognition, which lack frame-level discriminability and lead to fragmented pipelines.
This highlights the growing need for Frame-Discriminative frameworks that jointly learn spatial and temporal features from raw inputs. Such models promise better precision, adaptability, and generalization—particularly when built with task-specific backbones and unified, modality-aware training objectives.
4 Evaluation Metrics
In sports video event detection, multiple evaluation metrics are employed depending on the specific task—TAL, AS, or PES. These metrics capture different aspects of detection performance, including localization precision, recall quality, and robustness across varying thresholds.
4.1 Mean Average Precision (mAP)
For TAL tasks, the most widely used metric is Mean Average Precision (mAP), which measures how accurately models localize and classify actions. Given a ranked list of predicted intervals with associated confidence scores, precision and recall are defined as:
$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{\text{Total GT}},$ (1)
where $TP$ denotes true positives, $FP$ denotes false positives, and Total GT represents the number of ground-truth instances.
A prediction is considered a true positive if its Temporal Intersection over Union (T-IoU) with the ground truth exceeds a predefined threshold:
[…]
where $I_{p}$ and $I_{g}$ represent the predicted and ground-truth intervals, respectively.
The Average Precision (AP) for each class is calculated as:
$\mathrm{AP}=\sum_{k}(R_{k}-R_{k-1})P_{k},$ (3)
where $P_{k}$ and $R_{k}$ denote the precision and recall at the $k^{\text{th}}$ threshold. The mAP is obtained by averaging AP scores across all $C$ classes:
$\mathrm{mAP}=\frac{1}{C}\sum_{c=1}^{C}\mathrm{AP}_{c}.$ (4)
mAP offers a comprehensive evaluation by jointly considering localization and classification accuracy. It is widely accepted across different datasets, providing a common basis for benchmarking. However, it is sensitive to the choice of T-IoU thresholds (e.g., mAP@0.5 or mAP@0.75), and can penalize minor temporal misalignments heavily, which might be overly strict for fast or ambiguous actions, such as passes in soccer. Additionally, mAP alone does not separately assess the quality of action proposals independent of classification.
4.2 Average Recall at Average Number of Proposals (AR@AN)
In the context of TAPG, the Average Recall at Average Number of Proposals (AR@AN) is commonly employed. This metric evaluates the percentage of ground-truth instances that are successfully covered by at least one proposal, given a fixed number of proposals per video. Formally, recall is defined as:
Recall=Number of matched ground-truth instancesTotal number of ground-truth $\text{Recall}=\frac{\text{Number of matched ground-truth instances}}{\text{Total number of ground-truth instances}},$ (5)
where a match is determined based on a predefined T-IoU threshold.
AR@AN is typically reported at multiple proposal counts, such as AR@50, AR@100, or AR@200. It focuses on the completeness of the generated proposals, irrespective of the classification stage. However, it does not penalize redundancy or imprecise proposal boundaries, and therefore may overestimate performance in models that produce large or overlapping proposals.
4.3 Area Under the AR-AN Curve (AUC)
To assess proposal quality across a range of proposal counts, the Area Under the Curve (AUC) for the Average Recall versus Average Number of Proposals plot is calculated. Mathematically, it is defined as:
[…]
where $\text{AR}(n)$ denotes the average recall at $n$ proposals and $N$ is the maximum number of proposals considered.
AUC provides a holistic measure that captures how well a model maintains high recall across varying levels of proposal density. Models with higher AUC are generally more efficient, achieving strong recall with fewer proposals. Nevertheless, AUC can obscure detailed behaviors at specific proposal counts, which may be important in practical applications where the number of allowed proposals is constrained.
4.4 Tolerance Windows and AP Calculation in AS and PES
In contrast to TAL, which evaluates temporal intervals, both AS and PES treat detections as successful if the predicted timestamp falls within a predefined tolerance window centered around the ground-truth event. In AS, this window is typically broad, often ranging from 5 to 60 seconds, reflecting the nature of rapid, ambiguous actions where precise boundary annotation is difficult. PES, on the other hand, imposes a much stricter tolerance, typically constrained to 0–2 frames, requiring frame-level temporal precision.
Given this setup, the evaluation of AS and PES differs slightly from classical TAL evaluation. Rather than computing mAP over temporal intervals, AS and PES typically compute AP based on spot predictions relative to event anchors. Specifically, detections are first sorted by descending confidence scores. A prediction is considered a true positive if it falls within the defined tolerance window of any unmatched ground-truth event. Once a ground-truth event has been matched, it cannot be matched again.
For each true positive, the precision is computed at the time the true positive is found:
$\text{Precision}_{n}=\frac{\text{Number of True Positives up to }n}{\text{Number of Predictions up to }n}.$ (7)
The final AP is calculated as the average of the precision values at each true positive detection:
AP=1Total Number of Ground Truth $\mathrm{AP}=\frac{1}{\text{Total Number of Ground Truth Events}}\sum_{i=1}^{N}\text{Precision}_{i},$ (8)
where $N$ is the number of true positive detections.
This AP computation style ensures that early, high-confidence correct predictions contribute more to the final AP score, while late or incorrect predictions penalize the metric through lower precision values.
In AS, where larger tolerance windows are allowed, this formulation reflects the need for robust, high-recall detection under temporal ambiguity. However, the leniency of the large window means that models can achieve high AP even if the predicted timestamp is not highly precise. In contrast, PES evaluation, with its tight tolerance windows, demands much greater temporal accuracy; even small errors in frame prediction can result in a miss, making the task significantly harder and the AP score much more sensitive to prediction precision.
Thus, while the AP formulation for AS and PES follows a similar procedure, the strictness of the tolerance window fundamentally shifts the difficulty and interpretation of the resulting scores.
A key limitation of current AS/PES evaluation is its sensitivity to low confidence thresholds and multi-label predictions. Allowing multiple predictions per frame (e.g., with a 0.1 threshold) increases the chance of hitting the tolerance window and boosting AP, even if other incorrect predictions don’t contribute false positives—resulting in inflated mAP scores. This favors over-predictive models and may obscure true discriminative performance, especially in datasets with sparse or temporally isolated events.
To address this, future benchmarks should adopt stricter per-frame constraints—such as top-1 filtering, calibrated thresholds, or penalties for multi-class outputs—to ensure scores better reflect practical spotting accuracy.
While AS and PES share similar AP computation, the tolerance window and prediction filtering choices play a critical role in shaping fairness and interpretability.
5 Datasets
Datasets play a critical role in supervised deep learning, providing the foundation for both model training and evaluation. Transformer-based architectures [71, 52, 72] are particularly data-dependent, often requiring large-scale, high-quality datasets to achieve strong generalization. However, annotating sports videos remains a time-intensive and expertise-driven process. For example, distinguishing between different serve types in table tennis or tennis can be highly challenging due to subtle motion variations and the high speed of play [6]. Consequently, high-quality, precisely annotated datasets are especially valuable, as sports actions often exhibit limited generalization across different contexts.
In this section, we review publicly available sports-related datasets commonly used for event detection. It is also important to note that although some datasets are originally annotated with temporal intervals, they can often be adapted for AS and PES tasks by treating the start or end frame of each interval as a discrete event label. An overview of these datasets is presented in Table 3.
Table 3: Overview of sports-related datasets for event detection. Spotting denotes precise frame-level annotations, while Interval annotations specify action start and end times.
Dataset Sport Year Size / Duration Annotation Type Categories / Events
SoccerNet [20] Soccer 2018 500 videos / 764 hrs Spotting 3 (goals, cards, substitutions)
SSET [73] Soccer 2020 350 videos / 282 hrs Interval 11 event types, 15 story types
SoccerDB [74] Soccer 2020 346 videos / 669 hrs Interval 10
SoccerNet-v2 [57] Soccer 2021 500 videos / 764 hrs Spotting 17 event classes
SoccerNet Ball AS [55] Soccer 2023 7 videos Spotting 12 ball-action events
Tenniset [75] Tennis 2017 5 videos Interval 6
Tennis [47] Tennis 2022 3,345 clips Spotting 6 (court-specific ball contacts)
OpenTTGames [76] Table Tennis 2020 12 videos Spotting 3
P²A [77] Table Tennis 2024 2,721 videos / 272 hrs Interval 14 fine-grained / 8 high-level stroke classes
NCAA [78] Basketball 2016 257 videos /  1.5 hrs each Interval 14 (e.g., 3-point, dunk, steal)
Badminton Olympic [79] Badminton 2018 27 videos Interval 12
Figure Skating [22] Figure Skating 2021 11 videos Interval 4 transitions
FineDiving [80] Diving 2022 300 videos Interval 52 key pose transitions
FineGym [81] Gymnastics 2020 5,374 videos Spotting 32 fine-grained actions
MCFS [82] Figure Skating 2021 11,656 segments / 17.3 hrs Interval 130 across 4 event sets
5.0.1 Soccer
SoccerNet [20] is a large-scale dataset for soccer action localization, comprising 500 full-match videos from major European championships between 2015 and 2017. The dataset spans 764 hours of footage and occupies nearly 4TB of storage. Most event annotations are obtained by parsing official match reports. SoccerNet supports both AS and PES tasks, with manual one-second resolution annotations provided for three event types. For example, the ”card” event marks the moment a referee issues a yellow or red card due to a foul or misconduct.
SSET [73] is a smaller dataset relative to SoccerNet, containing 350 videos covering multiple soccer games, totaling 282 hours. It defines 11 event types and 15 story types. Designed primarily for TAL, its event annotations are interval-based. For instance, a ”kick” event is annotated from the moment a key player prepares to kick until the ball lands or exits the field. With slight annotation modifications, SSET can also be adapted for AS and PES.
SoccerDB [74] contains 346 high-quality soccer match videos, incorporating 270 matches from SoccerNet and 76 matches from the Chinese Super League (2017–2018) and FIFA World Cup editions. The dataset occupies 1.4TB and has a total duration of 668.6 hours. It defines 10 soccer event types with clear temporal boundaries, making it highly suitable for event detection tasks.
SoccerNet-v2 [57] builds upon SoccerNet by expanding the number of action classes from 3 to 17. Newly annotated events include ”Ball out of play,” ”Throw-in,” ”Foul,” ”Indirect free kick,” ”Clearance,” and ”Shot on target,” among others. Additionally, each timestamp is assigned a visibility tag, indicating whether the action is explicitly shown in the broadcast or inferred from gameplay context. For instance, a shot off target that is not directly shown but inferred based on subsequent play dynamics presents a unique challenge for automated detection.
SoccerNet Ball Action Spotting [55] introduces a new task focused on fine-grained ball interactions. Unlike the original SoccerNet events, which mainly targeted major game-changing actions, this task requires precise localization of frequent ball-related events, offering finer temporal granularity. The initial release included two key action classes: ”pass” (the moment the ball leaves a player’s foot) and ”drive” (the moment it is received by another player). The 2023 release provided annotations for 7 full matches, totaling 11,041 labeled timestamps. In 2024, the dataset was further expanded to include 12 distinct ball action classes, reflecting growing interest in modeling the flow of gameplay through continuous ball interactions.
5.0.2 Tennis
Tenniset [75] consists of five full-match videos from the 2012 London Olympic Games, sourced from YouTube. It defines six event types, such as ”set,” ”hit,” and ”serve,” annotated with precise temporal intervals. In addition, Tenniset provides textual descriptions of actions, such as ”quick serve is an ace,” enabling multimodal learning that combines video and text modalities.
The Tennis dataset [47], built upon the Vid2Player dataset [21], comprises 3,345 video clips from 28 professional tennis matches recorded at 25 or 30 FPS. Events are categorized into six classes, including ”player serve ball contact,” ”regular swing ball contact,” and ”ball bounce,” further divided based on court side (near or far court). The dataset supports fine-grained action spotting in tennis and facilitates evaluation under strict temporal precision settings.
5.0.3 Table Tennis
OpenTTGames [76] consists of 12 high-definition table tennis match videos recorded at 120 FPS, containing 4,271 labeled events. The dataset includes three event types: ball bounces, net hits, and empty events, all annotated with frame-level precision. OpenTTGames is particularly suited for training models targeting bounce event detection under high-speed gameplay conditions.
P2A [77] is one of the largest publicly available table tennis datasets, comprising 2,721 untrimmed broadcast videos with a total length of 272 hours. The dataset features footage from major international and Chinese tournaments, including the Tokyo 2020 Summer Olympics, World Cups, and World Championships. Annotations were conducted in collaboration with professional players and referees, defining 14 fine-grained stroke classes, which are further grouped into 8 higher-level action categories. Each action is annotated with precise start and end frames, enabling accurate event localization. The reliability of the dataset is strengthened by professional validation of the annotations.
5.0.4 Other Sports
NCAA [78] consists of 257 untrimmed college basketball game videos, each approximately 1.5 hours long. The dataset provides 14,548 video segments, with precise start and end times for 14 action categories, supporting temporal action localization tasks.
Badminton Olympic [79] contains 27 badminton match videos sourced from the official Olympic YouTube channel. It includes time-interval annotations for 12 action types, such as ”serve,” ”backhand,” and ”smash,” as well as point-level annotations, making it suitable for both action spotting and match-level analysis.
BadmintonTrack [83] is another badminton dataset, comprising 77,000 annotated frames from 26 unique singles matches filmed from an overhead broadcast-view camera. Originally, timestamp information indicating when a player struck the shuttlecock was included [84], although this metadata has since been removed in updated versions.
The Figure Skating dataset [22] contains 11 videos recorded at 25 FPS, featuring 371 short program performances from the 2010–2018 Winter Olympics and the 2017–2019 World Championships. It defines four transition event types critical for evaluating temporal precision in figure skating analysis.
FineDiving [80] includes 300 professional diving videos sourced from major competitions such as the Olympics, World Cup, and World Championships. The dataset defines 52 fine-grained action types, along with 29 sub-action types and 23 difficulty levels, offering a comprehensive benchmark for procedural action quality assessment.
FineGym [81] provides 5,374 gymnastics performances from international competitions. Each video is annotated with a hierarchical structure categorizing 32 spotting classes, covering disciplines such as balance beam and floor exercise, enabling fine-grained action spotting and classification.
MCFS [82] is a large-scale figure skating dataset comprising 11,656 video segments across 38 competitions, totaling 17.3 hours and 1.7 million frames. Annotations follow a hierarchical structure of 4 high-level action sets, 22 subsets, and 130 element actions, making MCFS highly suitable for dense temporal action localization tasks in figure skating.
One core limitation of current datasets is that most are recorded from professionally broadcasted footage with high-quality cameras, ideal angles, and stable viewpoints. While valuable for research, this setting does not reflect the reality faced by many sports analysts and practitioners, especially at non-elite or amateur levels. In many cases, analysts must rely on self-recorded videos with limited camera coverage, occlusions, and inconsistent viewpoints. As a result, models trained solely on broadcast footage may struggle to generalize to practical deployment scenarios where recording conditions are far less controlled.
6 Challenges and Future Directions
In this section, we critically examine current challenges in sports event detection and outline specific, actionable future research directions to address these limitations.
6.1 Generalization Across Diverse Sports
While many AS and PES models achieve strong results within individual sports—particularly soccer, as showcased by SoccerNet—they often rely heavily on domain-specific visual and contextual cues such as camera angles, common action semantics, and gameplay structure. This reliance limits their transferability to sports with different visual dynamics, motion patterns, and temporal scales.
A core limitation in current methods is the dependence on backbone architectures originally developed for image classification or coarse-grained action recognition. These models typically process video in fixed-length chunks and apply spatial-temporal aggregation, which diminishes subtle frame-level distinctions that are essential for spotting tasks. In contrast, AS and PES require temporally fine-grained representations that can detect minimal variations between adjacent frames—such as a foot making contact with a ball, or a player entering the field.
To improve generalization and robustness, future work should prioritize the development of frame-level feature extractors tailored specifically to the demands of sports video spotting. Promising directions include:
  • Designing encoders that preserve local temporal granularity, using lightweight 1D CNNs, temporal contrastive learning, or frame-attentive modules to enhance discriminative capacity.
  • Leveraging multimodal pretraining (e.g., CLIP [85]) to align visual, textual, and audio cues into semantically rich frame embeddings suitable for cross-domain transfer.
  • Exploring adaptive frame sampling strategies that focus representational capacity on moments of high temporal importance, improving both efficiency and localization accuracy.
By enhancing frame-wise representation learning, future AS and PES models will be better equipped to operate across a broader range of sports scenarios, improving their temporal precision and reducing reliance on domain-specific heuristics.
6.2 Unsupervised and Low-Supervision Methods
Creating extensive labeled datasets for sports event detection is costly, labor-intensive, and often demands expert domain knowledge, particularly in technical sports such as gymnastics, tennis, and figure skating. To mitigate these barriers, recent methods have begun exploring low-supervision learning paradigms such as knowledge distillation and active learning [61, 68]. These approaches significantly reduce dependency on large annotated datasets by transferring knowledge from pretrained models or selectively labeling informative samples.
Nevertheless, fully unsupervised or self-supervised methods remain largely unexplored. Future research directions include:
  • Development of self-supervised frameworks that leverage temporal consistency, contrastive learning, and multimodal alignment to extract meaningful event representations from unlabeled or weakly labeled video data.
  • Integration of unsupervised learning with domain adaptation techniques to further boost model performance and generalization across multiple sports domains.
These strategies will be essential for building scalable, efficient, and broadly applicable event detection models suitable for diverse sports analytics.
6.3 Enhanced Multimodal Fusion Approaches
Although existing AS and PES methodologies predominantly rely on visual data, audio cues can significantly enrich the detection of critical events in sports. Examples include sounds of ball impacts, crowd reactions, or figure skaters landing on ice—all of which provide valuable complementary temporal signals.
Currently, multimodal models predominantly employ straightforward fusion methods such as concatenation or late fusion [58, 59], which do not fully capture the complex interactions between modalities. To address these limitations, future research should explore more sophisticated fusion techniques:
  • Attention-based cross-modal transformers and gated attention mechanisms, capable of dynamically weighting and integrating audio and visual modalities.
  • Modality-specific encoders coupled with temporal alignment mechanisms to capture precise event timings, especially in noisy or visually ambiguous scenarios.
  • Exploiting commentary audio through automatic speech recognition and natural language processing to provide additional semantic context and weak supervision signals, aligning spoken descriptions directly with visual event occurrences.
Advancing multimodal fusion methodologies represents a crucial opportunity to significantly enhance the accuracy, robustness, and practical applicability of AS and PES frameworks in sports video analytics.
6.4 Real-World Applications: Gaps in Datasets and Evaluation Protocols
Despite impressive progress in AS and PES research, a substantial gap remains between academic benchmarks and real-world deployment scenarios. Most existing datasets are curated from professionally broadcasted footage, captured using high-quality cameras, ideal lighting, and fixed angles. While this enables consistent annotations and controlled evaluations, it fails to reflect the challenges encountered by sports analysts, coaches, and practitioners operating outside elite or televised settings. In many cases, especially at the amateur or semi-professional level, analysts rely on self-recorded videos using handheld devices or static single-angle setups, often under suboptimal conditions. Models trained exclusively on curated datasets may struggle to generalize to such noisy, diverse input sources.
In addition to data quality limitations, current evaluation metrics also fall short of representing practical constraints. Many AS and PES benchmarks compute mean mAP using low confidence thresholds (e.g., 0.1), allowing multiple class predictions per frame. While this boosts recall and improves mAP by capturing more true positives, it introduces an inflated performance signal—particularly problematic in domains such as racket sports, where only one event (e.g., hit, bounce) occurs at a time. In such cases, multi-label predictions per frame are unrealistic and may even degrade performance in downstream applications like rule enforcement or tactical feedback.
To address these shortcomings, future research should aim to:
  • Develop and evaluate models on datasets recorded in unconstrained, real-world environments to ensure robustness beyond broadcast-quality footage.
  • Introduce evaluation protocols that penalize over-prediction and reward frame-level discriminability—e.g., top-1 class selection per frame, or class-wise calibrated thresholds that reflect deployment precision requirements.
  • Encourage the release of datasets collected across a wider range of venues, competition levels, and camera setups to bridge the domain gap between research and practice.
Closing these gaps between academic settings and real-world use cases is critical for building sports video analysis systems that are not only accurate in theory but also reliable, efficient, and trustworthy in practice.
7 Conclusion
In this survey, we presented a comprehensive review of the current state of deep learning-based video event detection, with a particular focus on the critical tasks of TAL, AS, and PES within the sports analytics domain. By systematically categorizing and evaluating recent datasets and advanced methodologies, we identified key limitations, including poor generalizability across diverse sports contexts, reliance on extensive annotated datasets, and the underutilization of multimodal information sources.
Our analysis underscores the pressing need for the development of frame-level, generalizable frameworks that can robustly operate across diverse sports domains. Achieving this will likely require advances in task-specific backbone design, as well as the cross-domain evaluation strategies to ensure adaptability and performance consistency in varied real-world scenarios. Moreover, advancing unsupervised and low-supervision approaches, such as self-supervised learning and active learning, will be critical to reducing annotation costs and enabling scalable deployment. In parallel, future research should explore sophisticated multimodal fusion techniques that effectively integrate visual, audio, and textual modalities to enhance temporal precision, robustness, and event understanding in complex real-world environments.
By addressing these challenges, future efforts can significantly advance the capabilities of automated sports video analytics, leading to more comprehensive, accurate, and efficient systems capable of delivering actionable insights across a wide range of sports applications.